OpenStudy (anonymous):

the density of a block is 3.50g/cm3. the mass of the object is 360.15g. one side of the block measures 3.5cm. another side of the block measures 4.2 cm . what is the measurement of the third side?

OpenStudy (matt101):

The density is the mass per unit volume. That means that if we divide the whole mass in this situation by the whole volume, we should get the density. The volume of a block is length x width x height: \[3.5 = \frac{ 360.15 }{ 3.5 \times 4.2 \times h }\] Solving for h, you should get 7 cm!

OpenStudy (anonymous):

Thank you!!